OpenStudy (anonymous):

Consider the function f(x) = x(squared) − 2x + 6. Find the value of f(x) when x = 3 and select the correct answer below. A)1 B)3 C)6 D)9

OpenStudy (anonymous):

@dawsonj

OpenStudy (kittiwitti1):

This is the Biology forum. Are you looking for Mathematics?\[f(x)=-2x+6\rightarrow f(3)=-2(3)+6\]

OpenStudy (dawsonj):

c

OpenStudy (kittiwitti1):

Woops, forgot the beginning part.

OpenStudy (anonymous):

my bad xD @kittiwitti1 it was an honest mistake but ty and ty @dawsonj

OpenStudy (dawsonj):

np

OpenStudy (kittiwitti1):

\[f(x)=x^{2}-2x+6\rightarrow f(3)=(3)^{2}-2(3)-6\]

OpenStudy (kittiwitti1):

Also, it's preferable to not give direct answers unless there is no possible way to explain things. Just an FYI

OpenStudy (kittiwitti1):

...I messed up the sign, it's +6 sorry lol\[f(3)=(3)^{2}-2(3)+6\rightarrow 9-6+6\]....Doesn't that make it 9??

OpenStudy (anonymous):

I really just need the direct answer right now because its been explained to me over and over and i still don't get it and i'm in a super duper time crunch right now.

OpenStudy (kittiwitti1):

Just replace x with 3 ☻

OpenStudy (anonymous):

ty (:

OpenStudy (kittiwitti1):

np ^-^ I'll be here if you get stuck on it :P

OpenStudy (anonymous):

need to have WAY to many assignments turned in by 11 tomorrow in WAY too many classes :( So i'm on a major crunch for time :/

OpenStudy (anonymous):

awe, thank you (:

OpenStudy (kittiwitti1):

Though, what I solved for got me 9 and not 6... the answer should be D. I think you should check it yourself though x_x

OpenStudy (kittiwitti1):

Ok, yes, the answer is 9. https://www.wolframalpha.com/input/?i=x2-2x%2B6+when+x%3D3

OpenStudy (anonymous):

I should check, i'm just rushing and clicking tho tbh

OpenStudy (anonymous):

ty (: you're very helpful !

OpenStudy (kittiwitti1):

xD you can always use wolfram alpha when you're stuck on functions ^-^

OpenStudy (kittiwitti1):

To check*

OpenStudy (anonymous):

OMG TYSM YOU ARE GREAT !!

OpenStudy (anonymous):

For the functions f(x) = 2x + 3 and g(x) = 6x + 2, which composition produces the greatest output? I typed that in there and it told me it didn't understand -.-

OpenStudy (kittiwitti1):

I think you just stick in some random integer for both equations and see which one gives a greater result. Output usually means "solve for f(x)", with x being the input.

OpenStudy (kittiwitti1):

So... something like x=7 (:

OpenStudy (kittiwitti1):

Like this: f(7)=2(7)+3 and g(7)=6(7)+2... which one gives the greater answer? :-)

OpenStudy (anonymous):

Ohkai, ty (:

OpenStudy (kittiwitti1):

np

OpenStudy (anonymous):

is x=4 a function ?

OpenStudy (anonymous):

@dawsonj

OpenStudy (dawsonj):

yes it s

OpenStudy (anonymous):

ty

OpenStudy (dawsonj):

not a function

OpenStudy (shansha1):

f(x)=x^2-2x+6 \\ \Downarrow \\
f(3)=3^2 -2 \times 3+6=9-6+6=9 The answer is D. 9.

OpenStudy (damonte212):

..